Two particles, one with constant velocity 50 ms and the other with uniform acceleration 10 ms, starts moving simultaneously from

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
Your question is incomplete.
However I have written here a complete question and it's answer.
 
2 particles one with constant velocity of 50m/s and the other with uniform acceleration of 10m/s2 start moving simultaneously from the same place in the same direction.They will be at a distance of 125 m from each other afte
 
Solution-
 
We can do it by relative motion concepts. From frame of 1st particle, u21 = -50 and a21 = 10.
s21 = 125 = u21 t + 1/2 a21 t2 = -50t + 1/2 10 t2 
=> 25 = -10t + t2
Solving: t = 5 + 5 sqrt(2) = 5 (1 + sqrt(2)) seconds
 
Regards
Arun (askIITians forum expert)